It has been found that on average, four motorbikes go through an N1 tollgate every 20 minutes. What is the probability that

No more than 2 motorbikes will go through the tollgate in any randomly selected 20-minute period? Interpret your answer.


1
Expert's answer
2021-10-24T15:53:48-0400

This is a Poisson distribution with the mean "\mu=4."

"P(X\le2)=P(X=0)+P(X=1)+P(X=2)=e^{-4}(\frac{4^0}{0!}+\frac{4^1}{1!}+\frac{4^2}{2!})="

"=0.0183+0.0733+0.1465=0.2381."